OpenStudy (amistre64):

i got calc2 finals in a few hours .... should i be nervous?

OpenStudy (amistre64):

i might need to refresh the sin cos integral stuff ..... just so its fresh

OpenStudy (anonymous):

Say "All iz well" while you are nervous... it will relax you.. this really works !! :)

OpenStudy (across):

Integrals of trigonometric functions are a pain to remember, but don't sweat it too much. :)

OpenStudy (amistre64):

\[\int sin^n(x)dx=-\frac{1}{n+1}sin^{n-1}(x)cos(x)+\frac{n}{n+1}\int sin^{n-2}(x)dx\] now to see where i forgot it

OpenStudy (mr.math):

I used to hate them too, but with practice you'll get used to them. And you shouldn't be nervous.

OpenStudy (mr.math):

This formula will not be in the exam :P

OpenStudy (amistre64):

denom is "n" and top is "n-1" ... sooo close lol

OpenStudy (ash2326):

amistre64 stop studying atleast an hour before test, don't even think about any formulae

OpenStudy (amistre64):

stop studying? i havent even started yet;)

OpenStudy (anonymous):

You shouldn't be nervous, but you probably will be anyway :-)

OpenStudy (amistre64):

dance class is the only one im really concerned with

OpenStudy (amistre64):

might be rough on the convergence tests, but i aint no Raabe

OpenStudy (ash2326):

then don't study in the last hour

OpenStudy (amistre64):

im going to go in there chanting pedmas and foil lol

OpenStudy (anonymous):

calc2? like me? I thought you were taking graduate math!

OpenStudy (across):

That's because he hangs around us a lot. ;)

OpenStudy (anonymous):

calc2 is easy don't sweat it too much :-D

OpenStudy (anonymous):

but I hang around here a lot and I am terrible at math :(

OpenStudy (mr.math):

@across, @agd and @amistre, I've created a new study group for Math problems that are more difficult than the average of the problems you can find in this group. The group is new and needs more members. If you're interested, join "Meta-math" :D

OpenStudy (across):

I joined it earlier today (TuringTest told me about it).

OpenStudy (mr.math):

That's great! :D